I am having a problem solving the following equation in terms of logarithms. Could someone please help?
Problem :
2^2x-4= 3^x-21

Answer by stanbon(75887)

I am having a problem solving the following equation in terms of logarithms. Could someone please help?
Problem :
2^(2x-4) = 3^(x-21)
Take the log of both sides:
(2x-4)log2 = (x-21)log3
2xlog2 - 4log2 = xlog3 - 21*log3
----
(2log2-log3)x = 4log(2) - 21*log(3)
0.12494x = -8.81543
x = -70.557281
Cheers,
Stan H.